OpenStudy (anonymous):

find the slope of the line passing through the points (-4,7),(-3,6)

OpenStudy (anonymous):

=6-7/-3 - (-4) =-1/1 =-1 m = -1

OpenStudy (anonymous):

y2 - y1 / x2 - x1